# Photon pair generation via down-conversion in III-V semiconductor microrings: modal dispersion and quasi-phase matching

### v1

@@ I Introduction
For microrings fabricated out of III-V semiconductors, a “built-in” quasi-phase matching condition is present due to the difference between the crystal’s natural reference frame and the local reference frame in which the fields propagate [9, 10]. By adequately designing the microring resonator, the quasi-phase matching condition can easily be satisfied, as the signal and idler effective refractive indices can have values similar to that of the effective refractive index of the pump due to modal dispersion [8, 7, 11, 12, 13]. Experiments where high rates of photon pairs were generated have been reported [8, 7], and different approaches to arrive at rate equations or conversion efficiencies for microrings have been presented [10, 14, 15, 16]. In older work losses were neglected [10, 16], and in more recent work, they have been treated by an approach based on Langevin equations [14, 15].
In this work we extend these earlier calculations. We present the generation efficiencies of photon pairs by SPDC in a lossy III-V semiconductor microring resonator point-coupled to a waveguide, where the waveguide is pumped by CW or pulsed excitation in regimes where pump depletion can be neglected. We account for loss by adding a point-coupled “phantom” waveguide. Our approach uses an expansion in terms of asymptotic-in and asymptotic-out fields, which can be generalized to more complicated structures. This provides us with the statistics of not only the pairs of generated photons leaving the waveguide, but also the statistics of the pairs where both the photons are scattered, and those of the pairs where one photon is scattered and the other leaves the waveguide [17, 18]. Since the information of the scattered photons is not lost, we retain a full characterization of the outgoing quantum state. In the text we treat the scenario where the signal and idler photons generated are associated with different ring resonances (non-degenerate SPDC), but in Appendix B we present the corresponding expressions for the scenario where the signal and idler photons are associated with the same ring resonance (degenerate SPDC).
For excitation by pump pulses we address two different regimes. In the “pair-regime,” the generated state consists approximately of vacuum and a small probability amplitude of a pair of photons that can be written from first-order perturbation theory as
@@
The structure of this paper is as follows. In Sec. II, we begin by presenting the system and the theory to obtain the rate equation for CW excitation in the pair-regime. We also present the BWF and the number of photons in the generated squeezed state for pulsed excitation for both the pair-regime and the LOS regime. In Sec. III, we present a detailed sample calculation of the generation efficiencies of a realistic, simple structure: a microring coupled to a single physical waveguide, both fully encapsulated in silicon oxide (SiO2) cladding. We show a number of structures for which conversion efficiencies are on the order of $10^{-5}$ at critical coupling for CW low pump powers, where the quasi-phase matching and the dispersion properties of the system are taken into account, and show how the different coupling regimes compare to the one of critical coupling. We present various BWFs for the optimized structures for pulsed excitation for various pulse durations. The outgoing states have a range of different Schmidt numbers ([19, 20, 21]). We also show that our optimized structures are potentially able to generate high amounts of squeezing and could possibly lead to pump depletion – which in turn could lead to deterministic generation of non-Gaussian states [22] – for relatively low energy pump pulses. Finally, in Sec. IV, we include a brief conclusion, and discuss future work.
@@ II Theory
We envision a pump pulse or CW excitation coupled from the waveguide into the ring at frequencies within one of the ring resonances, and photon pairs either generated at frequencies within different ring resonances (non-degenerate SPDC, considered in the text), or within one resonance (degenerate SPDC, considered in Appendix B).
@@ II.1 The linear Hamiltonian
where $\varepsilon_{1}(\bm{r}_{\perp},\omega_{u})$ is the relative permittivity, $v_{g}(\bm{r}_{\perp};\omega_{u})$ is the local group velocity of the material, and $v_{p}(\bm{r}_{\perp};\omega_{u})$ is the local phase velocity of the material, all at frequency $\omega_{u}$ [20]. Here we have used the fact that the dot product $\bm{\mathsf{d}}_{u}^{*}(\bm{r}_{\perp},\zeta)\cdot\bm{\mathsf{d}}_{u}(\bm{r}_{\perp},\zeta)$ is independent of $\zeta$, and we have written it simply as $\bm{\mathsf{d}}_{u}^{*}(\bm{r}_{\perp})\cdot\bm{\mathsf{d}}_{u}(\bm{r}_{\perp})$. In the systems treated here we will consider pump frequencies near a ring resonance frequency that we denote by $\omega_{P}$, signal frequencies near a ring resonance that we denote by $\omega_{S}$, and idler frequencies near a ring resonance that we denote by $\omega_{I}$; for each pump scenario we will also consider a range of different $\omega_{S}$ and $\omega_{I}$.
@@ II.2 Asymptotic fields
The mode fields $\bm{D}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ are constructed (see Fig. 2b and 2c) so that they are equal to $\bm{D}_{uk}(\bm{r})$ for waveguide $\lambda$ and $s>0$ (indicated by the red arrows), and vanish for the other waveguide for $s>0$; the mode fields $\bm{D}_{uk}^{\mathrm{in},\lambda}(\bm{r})$ (see Figure 2a for the asymptotic-in mode field for the actual waveguide) are constructed so that they are equal to $\bm{D}_{uk}(\bm{r})$ for waveguide $\lambda$ and $s<0$ (indicated by the red arrow) and vanish for the other waveguide for $s<0$. So the asymptotic-out expansion is appropriate for the signal and idler fields leaving the structure, while the asymptotic-in expansion is appropriate for the pump field entering the structure [18, 17, 23].
We will assume that the nonlinearity is effective only in the ring resonator, where the fields are enhanced. Thus, in constructing the nonlinear interaction Hamiltonian we will need the expressions for the fields $\bm{D}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ and $\bm{D}_{uk}^{\mathrm{in},\lambda}(\bm{r})$ in the ring. We denote these by $\bm{\mathsf{D}}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ and $\bm{\mathsf{D}}_{uk}^{\mathrm{in},\lambda}(\bm{r})$, and they are given by [17]:
@@ II.3 The nonlinear Hamiltonian
We note here that the primed indices $\{i^{\prime},j^{\prime},k^{\prime}\}$ refer to the lab frame, and the unprimed indices $\{i,j,k\}$ to the ring frame, as described in Appendix A and indicated in Fig. 1. The relevant term for the generation of a pair of non-degenerate signal and idler photons, inside frequency bins $S$ and $I$ (where we take $\omega_{S}>\omega_{I}$), from a pump photon inside the bin labelled $P$ is
$\begin{split}H^{\mathrm{SPDC}}_{NL}={-}\frac{2}{\epsilon_{0}}\int\Gamma^{(2)}_{i^{\prime}j^{\prime}k^{\prime}}(\bm{r})&\left[{\mathsf{D}}_{S}^{\mathrm{out}(i^{\prime})}(\bm{r}){\mathsf{D}}_{I}^{\mathrm{out}(j^{\prime})}(\bm{r})\right]^{\dagger}\\ &\times{\mathsf{D}}_{P}^{\mathrm{in}(k^{\prime})}(\bm{r})\mathrm{d}\bm{r}+\mathrm{H.c.}.\end{split}$ (32)
The pump photons enter the ring from the actual waveguide, and the generated signal and idler photons exit the ring either by the phantom or the actual waveguide, depending on the label $\lambda$ and $\lambda^{\prime}$. Using Eq. (23) in Eq. (32) we obtain
@@
which are the possible quasi-phase matching conditions. Further, at most one of $V_{SIP}^{(\pm)}$ can vanish; when the $+$ (--) quasi-phase matched condition is met, the value of $V_{SIP}^{(+)}$ ($V_{SIP}^{(-)}$) is equal to $2\pi\mathcal{R}$ ($2\pi\mathcal{R}$) and the $V_{SIP}^{(-)}$ ($V_{SIP}^{(+)}$) term is zero, and so at most one term in Eq. (38) contributes to $\bar{\mathsf{K}}_{SIP}$. We write the “$\pm$ matched” version of Eq. (35) by combining Eq. (38) and the quasi-phase matched $V_{SIP}^{(\pm)}$ terms,
@@ II.4.1 Rates
is the rate of generation of photon pairs where the signal photon ($S$) is leaving through the $\lambda$ waveguide and the idler photon ($I$) is leaving through the $\lambda^{\prime}$ waveguide. After integrating over both $k_{1}$ and $k_{2}$, we find
@@
We note that the rate expression from Eq. (54) is for a given set of resonances with resonant wavenumbers satisfying the quasi-phase matching condition given by Eq. (41); if this condition is not met, within our treatment of the coupling of the resonator to the channel the generation rate strictly vanishes. The value of the effective area that should be used in the rate equation (either $A_{\mathrm{eff}}^{(+)}$ or $A_{\mathrm{eff}}^{(-)}$) is chosen by which quasi-phase matching condition is met from Eq. (41) (either $+$ or --). We have written the rate as a function of the vacuum power [16, 17]:
@@
where $S$ and $I$ label the distinct frequency bins, with $S>I$ indicating that the centre frequency of the signal bin ($\omega_{S}$) is taken greater than the centre frequency of the idler bin ($\omega_{I}$). The expression for the non-degenerate rate $R^{SI}_{\mathrm{ac,ac}}$ follows from Eq. (53), and the degenerate efficiency $\hat{R}^{SS}_{\mathrm{ac,ac}}$ is that of Eq. (118). We can also introduce a generation efficiency $\mathcal{E}^{\mathrm{total}}_{\mathrm{ac,ac}}$, defined as the ratio of the rate at which photon pairs are generated in the actual channel to the rate at which pump photons are incident,
@@ II.4.2 Discussion
First, the rate is inversely proportional to the effective area, and so it scales with the square of the integral in Eq. (43). Maximizing this integral involves identifying the best overlap between the electric field mode profiles.
@@
Third, consider an $R^{SI}_{\lambda\lambda^{\prime}}$ that does not vanish; this arises when the quasi-phase matching condition from Eq. (42) is satisfied. If the pump frequency were set to the ring resonance frequency $\omega_{P}$, and we wanted the signal and idler generated at the ring resonance frequencies $\omega_{S}$ and $\omega_{I}$ respectively, then from energy conservation we would require $\omega_{P}-\omega_{S}-\omega_{I}=0$. In general this cannot be satisfied together with the quasi-phase matching condition, which is required to have $R^{SI}_{\lambda\lambda^{\prime}}\neq 0$. So the signal and the idler will not be generated exactly at the ring resonance frequencies, which will weaken the generation rate.
@@ II.5 Pulsed excitation
II.5.1 Pair-regime
We first look at the pair-regime, where the outgoing state approximately consists of only pairs of photons. In this case, the output ket can be written as [25, 17]
@@ II.5.1 Pair-regime
which is written in terms of the components of the biphoton wave function (BWF) $\varphi^{\mathrm{II}}_{\lambda\lambda^{\prime}}(k_{1},k_{2})$ associated with the signal photon leaving the $\lambda$ waveguide and the idler photon leaving the $\lambda^{\prime}$ waveguide,
@@
where $\lvert\psi\rangle$ is the state of the generated photons, which we write in the pair-regime as
@@
In this pair-regime, the probability $|\beta|^{2}$ that a pair of photons is generated is determined by using the expression from Eq. (73) for the components of the biphoton wave function in the normalization condition from Eq. (66). We find
@@
Dividing by the number of pump photons, the efficiency for the pair-regime, which does not vary within this regime, can be identified as
$\begin{split}\mathcal{E}_{\mathrm{pulse;pair-regime}}^{SI}&=\frac{|\beta|^{2}}{|\alpha|^{2}},\end{split}$ (78)
@@
$\begin{split}[\mathcal{E}_{\mathrm{pulse;pair-regime}}^{SI}]_{\lambda\lambda^{\prime}}&=\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{|\alpha|^{2}},\end{split}$ (79)
@@
$\begin{split}\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{|\beta_{\mu\mu^{\prime}}|^{2}}=\frac{[\mathcal{E}_{\mathrm{pulse;pair-regime}}^{SI}]_{\lambda\lambda^{\prime}}}{[\mathcal{E}_{\mathrm{pulse;pair-regime}}^{SI}]_{\mu\mu^{\prime}}}=\frac{\eta_{S}^{\lambda}\eta_{I}^{\lambda^{\prime}}}{\eta_{S}^{\mu}\eta_{I}^{\mu^{\prime}}}=\frac{\Gamma_{S}^{\lambda}\Gamma_{I}^{\lambda^{\prime}}}{\Gamma_{S}^{\mu}\Gamma_{I}^{\mu^{\prime}}}.\end{split}$ (80)

### v3

# Photon pair generation via down-conversion in III-V semiconductor microrings: modal dispersion and quasi-phase-matching

@@ I Introduction
For microrings fabricated out of III-V semiconductors, a “built-in” quasi-phase-matching condition is present due to the difference between the crystal’s natural reference frame and the local reference frame in which the fields propagate [9, 10]. In addition, III-V semi-conductors have very high values of $\chi^{(2)}$ compared to other materials commonly used for nonlinear interactions [11], leading to more efficient generation of quantum states. Also, recent advances in fabrication techniques have resulted in high-Q microring resonators useful for down-conversion from the visible to the telecommunication wavelengths have been developed [7, 12]. By adequately designing the microring resonator, the quasi-phase-matching condition can easily be satisfied, as the signal and idler effective refractive indexes can have values similar to that of the effective refractive index of the pump due to modal dispersion [7, 8, 12, 13, 14].
Experiments where high rates of photon pairs were generated have been reported [7, 8], and different approaches to arrive at rate equations or conversion efficiencies for microrings have been presented [10, 15, 16, 17]. In older work losses were neglected [10, 17, 18], and in more recent work, they have been treated by an approach based on Langevin equations [15, 16]. In this work we extend these earlier calculations by treating the SPDC interaction shown in [18] with the strategy for lossy structures presented in [19]. We also present extensive numerical simulations of realistic microring resonators from III-V semiconductors for down-conversion from the visible to the telecommunication range. Our calculation also extends beyond the pair regime, providing efficiencies for the squeezing regime.
We present the generation efficiencies of photon pairs by SPDC in a lossy III-V semiconductor microring resonator point-coupled to a waveguide, where the waveguide is pumped by cw or pulsed excitation in regimes where pump depletion can be neglected. We account for loss by adding a point-coupled “phantom” waveguide. Our approach uses an expansion in terms of asymptotic-in and asymptotic-out fields, which can be generalized to more complicated structures. This provides us with the statistics of not only the pairs of generated photons exiting the waveguide, but also the statistics of the pairs where both the photons are scattered, and those of the pairs where one photon is scattered and the other leaves the waveguide [19, 18]. Since the information of the scattered photons is not lost, we retain a full characterization of the outgoing quantum state. The results from our approach can also be used directly to treat more sophisticated problems involving high squeezing, pump-depletion, and non-Gaussian state generation [20].
In the text we treat the scenario where the signal and idler photons generated are associated with different ring resonances (nondegenerate SPDC), but in Appendix B we present the corresponding expressions for the scenario where the signal and idler photons are associated with the same ring resonance (degenerate SPDC).
For excitation by pump pulses we address two different regimes. In the “pair regime,” the generated state consists approximately of vacuum and a small probability amplitude of a pair of photons that can be written from first-order perturbation theory as
@@
The structure of this paper is as follows. In Sec. II, we begin by presenting the system and the theory to obtain the rate equation for cw excitation in the pair regime. We also present the BWF and the number of photons in the generated squeezed state for pulsed excitation for both the pair regime and the LOS regime. In Sec. III, we present a detailed sample calculation of the generation efficiencies of a realistic, simple structure: a microring coupled to a single physical waveguide, both fully encapsulated in silicon oxide (SiO2) cladding. We show a number of structures for which conversion efficiencies are on the order of $10^{-5}$ at critical coupling for cw low pump powers, where the quasi-phase-matching and the dispersion properties of the system are taken into account, and show how the different coupling regimes compare with the one of critical coupling. We present various BWFs for the optimized structures for pulsed excitation for various pulse durations. The outgoing states have a range of different Schmidt numbers ([21, 22, 23]). We also show that our optimized structures are potentially able to generate high amounts of squeezing and could possibly lead to pump depletion – which in turn could lead to deterministic generation of non-Gaussian states [24] – for relatively-low-energy pump pulses. Finally, in Sec. IV, we include a brief conclusion, and discuss future work.
@@ II Theory
We envision a pump pulse or cw excitation coupled from the waveguide into the ring at frequencies within one of the ring resonances, and photon pairs either generated at frequencies within different ring resonances (nondegenerate SPDC, considered in the text), or within one resonance (degenerate SPDC, considered in Appendix B).
@@ II.1 The linear Hamiltonian
where $\varepsilon_{1}(\bm{r}_{\perp},\omega_{u})$ is the relative permittivity, $v_{g}(\bm{r}_{\perp};\omega_{u})$ is the local group velocity of the material, and $v_{p}(\bm{r}_{\perp};\omega_{u})$ is the local phase velocity of the material, all at frequency $\omega_{u}$ [22]. Here we have used the fact that the dot product $\bm{\mathsf{d}}_{u}^{*}(\bm{r}_{\perp},\zeta)\cdot\bm{\mathsf{d}}_{u}(\bm{r}_{\perp},\zeta)$ is independent of $\zeta$, and we have written it simply as $\bm{\mathsf{d}}_{u}^{*}(\bm{r}_{\perp})\cdot\bm{\mathsf{d}}_{u}(\bm{r}_{\perp})$. In the systems treated here we consider pump frequencies near a ring resonance frequency that we denote by $\omega_{P}$, signal frequencies near a ring resonance that we denote by $\omega_{S}$, and idler frequencies near a ring resonance that we denote by $\omega_{I}$; for each pump scenario we also consider a range of different $\omega_{S}$ and $\omega_{I}$.
@@ II.2 Asymptotic fields
The mode fields $\bm{D}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ are constructed (see Fig. 2b and 2c) so that they are equal to $\bm{D}_{uk}(\bm{r})$ for waveguide $\lambda$ and $s>0$ (indicated by the red arrows), and vanish for the other waveguide for $s>0$; the mode fields $\bm{D}_{uk}^{\mathrm{in},\lambda}(\bm{r})$ (see Figure 2a for the asymptotic-in mode field for the actual waveguide) are constructed so that they are equal to $\bm{D}_{uk}(\bm{r})$ for waveguide $\lambda$ and $s<0$ (indicated by the red arrow) and vanish for the other waveguide for $s<0$. So the asymptotic-out expansion is appropriate for the signal and idler fields exiting the structure, while the asymptotic-in expansion is appropriate for the pump field entering the structure [18, 19, 25].
We assume that the nonlinearity is effective only in the ring resonator, where the fields are enhanced. Thus, in constructing the nonlinear interaction Hamiltonian we need the expressions for the fields $\bm{D}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ and $\bm{D}_{uk}^{\mathrm{in},\lambda}(\bm{r})$ in the ring. We denote these by $\bm{\mathsf{D}}_{uk}^{\mathrm{out},\lambda}(\bm{r})$ and $\bm{\mathsf{D}}_{uk}^{\mathrm{in},\lambda}(\bm{r})$, and they are given by [19]:
@@ II.3 The nonlinear Hamiltonian
We note here that the primed indices $\{i^{\prime},j^{\prime},k^{\prime}\}$ refer to the laboratoryframe, and the unprimed indices $\{i,j,k\}$ to the ring frame, as described in Appendix A and indicated in Fig. 1. The relevant term for the generation of a pair of nondegenerate signal and idler photons, inside frequency bins $S$ and $I$ (where we take $\omega_{S}>\omega_{I}$), from a pump photon inside the bin labelled $P$ is
$\begin{split}H^{\mathrm{SPDC}}_{NL}={-}\frac{2}{\epsilon_{0}}\int\Gamma^{(2)}_{i^{\prime}j^{\prime}k^{\prime}}(\bm{r})&\left[{\mathsf{D}}_{S}^{\mathrm{out}(i^{\prime})}(\bm{r}){\mathsf{D}}_{I}^{\mathrm{out}(j^{\prime})}(\bm{r})\right]^{\dagger}\\ &\times{\mathsf{D}}_{P}^{\mathrm{in}(k^{\prime})}(\bm{r})\mathrm{d}\bm{r}+\mathrm{H.c.},\end{split}$ (32)
where the displacement operators $\bm{\mathsf{D}}_{u}^{\mathrm{in/out}}(\bm{r})$ are those from Eq. (22), and we have used the expression for the displacement fields $\bm{\mathsf{D}}^{\mathrm{in/out},\lambda}_{uk}(\bm{r})$ inside the ring from Eq. (23). The pump photons enter the ring from the actual waveguide, and the generated signal and idler photons exit the ring either by the phantom or the actual waveguide, depending on the label $\lambda$ and $\lambda^{\prime}$. Using Eq. (23) in Eq. (32) we obtain
@@
which are the possible quasi-phase-matching conditions. Further, at most one of $V_{SIP}^{(\pm)}$ can vanish; when the $+$ (--) quasi-phase-matched condition is met, the value of $V_{SIP}^{(+)}$ ($V_{SIP}^{(-)}$) is equal to $2\pi\mathcal{R}$ ($2\pi\mathcal{R}$) and the $V_{SIP}^{(-)}$ ($V_{SIP}^{(+)}$) term is zero, and so at most one term in Eq. (38) contributes to $\bar{\mathsf{K}}_{SIP}$. We write the “$\pm$ matched” version of Eq. (35) by combining Eq. (38) and the quasi-phase-matched $V_{SIP}^{(\pm)}$ terms,
@@ II.4.1 Rates
is the rate of generation of photon pairs where the signal photon ($S$) is exiting through the $\lambda$ waveguide and the idler photon ($I$) is exiting through the $\lambda^{\prime}$ waveguide. After integrating over both $k_{1}$ and $k_{2}$, we find
@@
We note that the rate expression from Eq. (54) is for a given set of resonances with resonant wavenumbers satisfying the quasi-phase-matching condition given by Eq. (41); if this condition is not met, within our treatment of the coupling of the resonator to the channel the generation rate strictly vanishes. The value of the effective area that should be used in the rate equation (either $A_{\mathrm{eff}}^{(+)}$ or $A_{\mathrm{eff}}^{(-)}$) is chosen by which quasi-phase-matching condition is met from Eq. (41) (either $+$ or --). We have written the rate as a function of the vacuum power [17, 19]:
@@
where $S$ and $I$ label the distinct frequency bins, with $S>I$ indicating that the centre frequency of the signal bin ($\omega_{S}$) is taken greater than the centre frequency of the idler bin ($\omega_{I}$). The expression for the nondegenerate rate $R^{SI}_{\mathrm{ac,ac}}$ follows from Eq. (53), and the degenerate efficiency $\hat{R}^{SS}_{\mathrm{ac,ac}}$ is that of Eq. (118). We can also introduce a generation efficiency $\mathcal{E}^{\mathrm{total}}_{\mathrm{ac,ac}}$, defined as the ratio of the rate at which photon pairs are generated in the actual channel to the rate at which pump photons are incident,
@@ II.4.2 Discussion
First, the rate is inversely proportional to the effective area, and so it scales with the square of the integral in Eq. (43). Maximizing this integral involves identifying the best overlap between the electric-field mode profiles.
@@
Third, consider an $R^{SI}_{\lambda\lambda^{\prime}}$ that does not vanish; this arises when the quasi-phase-matching condition from Eq. (42) is satisfied. If the pump frequency were set to the ring resonance frequency $\omega_{P}$, and we wanted the signal and idler generated at the ring resonance frequencies $\omega_{S}$ and $\omega_{I}$ respectively, then from energy conservation we would require $\omega_{P}-\omega_{S}-\omega_{I}=0$. In general this cannot be satisfied together with the quasi-phase-matching condition, which is required to have $R^{SI}_{\lambda\lambda^{\prime}}\neq 0$. So the signal and the idler will not be generated exactly at the ring resonance frequencies, which will weaken the generation rate.
@@ II.5 Pulsed excitation
II.5.1 pair regime
We first look at the pair regime, where the outgoing state approximately consists of only pairs of photons. In this case, the output ket can be written as [27, 19]
@@
which is written in terms of the components of the biphoton wave function (BWF) $\varphi^{\mathrm{II}}_{\lambda\lambda^{\prime}}(k_{1},k_{2})$ associated with the signal photon exiting the $\lambda$ waveguide and the idler photon exiting the $\lambda^{\prime}$ waveguide,
@@
where $\lvert\psi\rangle$ is the state of the generated photons, which we write in the pair regime as
@@
In this pair regime, the probability $|\beta|^{2}$ that a pair of photons is generated is determined by using the expression from Eq. (73) for the components of the biphoton wave function in the normalization condition from Eq. (66). We find
@@
Dividing by the number of pump photons, the efficiency for the pair regime, which does not vary within this regime, can be identified as
$\begin{split}\mathcal{E}_{\mathrm{pulse;pairregime}}^{SI}&=\frac{|\beta|^{2}}{|\alpha|^{2}},\end{split}$ (78)
@@
$\begin{split}[\mathcal{E}_{\mathrm{pulse;pairregime}}^{SI}]_{\lambda\lambda^{\prime}}&=\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{|\alpha|^{2}},\end{split}$ (79)
@@
$\begin{split}\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{|\beta_{\mu\mu^{\prime}}|^{2}}=\frac{[\mathcal{E}_{\mathrm{pulse;pairregime}}^{SI}]_{\lambda\lambda^{\prime}}}{[\mathcal{E}_{\mathrm{pulse;pairregime}}^{SI}]_{\mu\mu^{\prime}}}=\frac{\eta_{S}^{\lambda}\eta_{I}^{\lambda^{\prime}}}{\eta_{S}^{\mu}\eta_{I}^{\mu^{\prime}}}=\frac{\Gamma_{S}^{\lambda}\Gamma_{I}^{\lambda^{\prime}}}{\Gamma_{S}^{\mu}\Gamma_{I}^{\mu^{\prime}}}.\end{split}$ (80)